OpenStudy (anonymous):

Ok i'm having problems factoring 3x^2-4x+2

OpenStudy (anonymous):

\[3x^2-4x+2\]

OpenStudy (anonymous):

use the quadratic equation?

OpenStudy (dumbcow):

3*2 = 6 --> no factors of 6 add up to -4 Does not factor

OpenStudy (anonymous):

alright i guess i will use the quadratic, I just wanted to make sure their wasn't another pattern i was forgetting.

myininaya (myininaya):

\[b=(\frac{-4}{2}+z)+(\frac{-4}{2}-z)\] \[3(2)=6=ac=(\frac{-4}{2}+z)(\frac{-4}{2}-z)\] we need to solve \[6=(\frac{-4}{2}+z)(\frac{-4}{2}-z)=(\frac{-4}{2})^2-z^2\] so we have \[6=4-z^2=> z^2=4-6=-2 => z^2=-2 => z=\pm \sqrt{-2}=\pm i \sqrt{2}\] we will just be needing \[z=i \sqrt{2}\] so bx=\[(-2+i \sqrt{2})x+(-2-i \sqrt{2})x\] so we replace the bx term in ax^2+bx+c with the above \[3x^2+(-2+i \sqrt{2})x+(-2- i \sqrt{2})x+2\] now factor by grouping! \[3x(x+\frac{-2+i \sqrt{2}}{3})+(-2 - i \sqrt{2})(x+\frac{2}{-2- i \sqrt{2}})\] now we want to see if \[\frac{-2 + i \sqrt{2}}{3} and \frac{2}{-2- i \sqrt{2}}\] to do this we should rationalize the denominator of this second one \[\frac{2}{-2-i \sqrt{2}} \cdot \frac{-2+i \sqrt{2}}{-2+i \sqrt{2}}=\frac{2(-2+i \sqrt{2})}{4-i^2 (2)}=\frac{2(-2+i \sqrt{2})}{4-(-1)(2)}\] \[=\frac{2(-2+i \sqrt{2})}{6}=\frac{-2+ i \sqrt{2}}{3}\] so we have \[3x(x+\frac{-2+i \sqrt{2}}{3})+(-2 - i \sqrt{2})(x+\frac{-2+i \sqrt{2}}{3})\] so each term has the factor \[x+\frac{-2 + i \sqrt{2}}{3}\] in common we can factor that out \[(x+\frac{-2+i \sqrt{2}}{3})(3x-2-i \sqrt{2})\]

OpenStudy (anonymous):

This does not have real solutions ... its solutions are complex

OpenStudy (anonymous):

these being the solutions x = 1/3 (2-i sqrt(2)) x = 1/3 (2+i sqrt(2))

OpenStudy (anonymous):

Depending what level of math you are at. It cannot be factored unless you use imaginary numbers and you might not of be at that level yet. So it cannot be factored

OpenStudy (anonymous):

I'm getting \[\frac{4\pm2i \sqrt2}{6}\] can you show me where i'm going wrong?

myininaya (myininaya):

you can reduce but looks good! :)

OpenStudy (dumbcow):

that is correct, just needs to be simplified, factor out a 2

OpenStudy (anonymous):

oh doh! thanks guys